# Recurrent Self-Attention Dynamics: An Energy-Agnostic Perspective from Jacobians

### v2

@@ Abstract
The theoretical understanding of self-attention (SA) has been steadily progressing. A prominent line of work studies a class of SA layers that admit an energy function decreased by state updates. While it provides valuable insights into inherent biases in signal propagation, it often relies on idealized assumptions or additional constraints not necessarily present in standard SA. Thus, to broaden our understanding, this work aims to relax these energy constraints and provide an energy-agnostic characterization of inference dynamics by dynamical systems analysis. In more detail, we first consider relaxing the symmetry and single-head constraints traditionally required in energy-based formulations. Next, to investigate more general SA architectures capable of oscillatory dynamics without necessarily admitting an energy function, we analyze the Jacobian matrix of the state. We reveal that normalization layers effectively normalize the Jacobian’s complex eigenvalues, forcing the dynamics close to a critical state. This significantly enhances inference performance. Furthermore, we utilize the Jacobian perspective to develop regularization methods for training and a pseudo-energy for monitoring inference dynamics.
@@ 1 Introduction
The theoretical understanding of self-attention (SA), a central component of Transformer architectures (Vaswani, 2017), has deepened in recent years, including phenomena such as rank collapse (Noci et al., 2022) and expressive capacity (Yun et al., 2020; Kajitsuka and Sato, 2024; Kingma and Ba, 2025). One major line of research formulates attention mechanisms as processes that minimize explicit or implicit energy functions (Ramsauer et al., 2021; Yang et al., 2022; Hoover et al., 2023). Since these energy functions serve as potential functions for gradient flows or as Lyapunov functions, they offer convergence guarantees and provide intuitive explanations for behaviors such as clustering and rank collapse in recurrent SA architectures (Geshkovski et al., 2023a, b, 2024; Bruno et al., 2025). They restrict attention dynamics to a hypersphere, typically as a result of normalization. This facilitates theoretical analysis and can yield well-behaved functional properties (Castin et al., 2024).
@@
  • First, we revisit the energy-based formulation and partially relax traditional architectural constraints, such as symmetric weights and single-head assumptions, to better approximate realistic SA settings (Section 4). These relaxed constraints provide insights into designing regularization methods, which we experimentally explore later in Section 6.2.
  • To study a broader class of SAs more flexibly, we next analyze the Jacobian matrix of SA dynamics (Section 5). The Jacobian approach is more general than the energy-based analysis (a.k.a. Lyapunov’s direct method) in the sense that it characterizes linear stability (a.k.a. Lyapunov’s indirect method) and enables us to detect non-stationary dynamics, including oscillations. We demonstrate that normalization layers, unique to discrete updates, play a critical role in stabilizing dynamics. Specifically, they effectively suppress the Jacobian’s spectral norm (Proposition 5.1) and control oscillatory behaviors by normalizing the complex eigenvalues of the Jacobian (Section 5.2). Empirically, we confirm that high-performance SA models exhibit a maximum Lyapunov exponent close to zero, suggesting that rich non-stationary inference dynamics emerge at the boundary between convergence and instability.
  • Finally, we investigate test-time scaling of inference through the lens of Jacobians. We show that regularizing the spectral norm of weight matrices in SA improves performance (Section 6.2), and that the Jacobian offers an interpretation of the pseudo-energy proposed in prior work, linking it to large eigenvalues (Section 6.3).
Thus, our work broadens the dynamical understanding of SA through dynamical systems analysis, laying a foundation for further exploration of realistic SA architectures.
@@ 2 Related work
Geshkovski et al. (2023a, b, 2024); Karagodin et al. (2024); Bruno et al. (2025) formulated recurrent SA dynamics as interactions among tokens (“particles”), enabling theoretical analysis of phenomena such as meta-stable clustering and rank-one collapse. Their continuous-time dynamics monotonically decrease an energy (Lyapunov) function, typically requiring constraints like single-head attention, hyperspherical token states, and symmetric weights. Yang et al. (2022) similarly interpreted Transformers as alternating minimization of energy functions, though with stricter conditions on step sizes and fixed-point proximity. Ramsauer et al. (2021) formalized the attention mechanism as modern Hopfield networks, and Hoover et al. (2023); Hu et al. (2025) further developed energy functions for Transformers. We do not address approaches based on the Hopfield networks, as they require architectural modifications distinct from standard self-attention.
@@ 3 Preliminaries
Geshkovski et al. (2023a); Karagodin et al. (2024) used continuous equations and particle interpretation of tokens to model state-update dynamics of SA as:
@@
depending on the analyses. (Bruno et al., 2025) further assumes an unnormalized version of softmax. Under these conditions, the SA update can decrease an energy function. That is, the dynamics evolve in a way that monotonically decreases the energy, thereby ensuring the Lyapunov stability. Because these models suppose symmetric weights, we refer to a class of SA layers with symmetric weights and Lyapunov functions as symmetric SA. Yang et al. (2022) also formalized updates of SA using a symmetric matrix  (Appendix C).
@@
To facilitate theoretical analysis of SA, several studies (Geshkovski et al., 2023a; Miyato et al., 2025) have introduced a spherical constraint on token vectors by enforcing by requiring that each token vector has unit norm. This constraint enables an interpretation of token interactions as dynamics of particles on a hypersphere, and plays a key role in controlling the Lipschitz continuity of SA (Castin et al., 2024). There are two commonly used operators with a spherical constraint: a normalization operator $\Pi$ that enforces the spherical constraint, and a projection operator $\operatorname{Proj}$ that projects onto the tangent space of the sphere. Given a token matrix $\bm{X},\bm{Y}\in\mathbb{R}^{S\times D}$ such that $\|\bm{X}_{[i,:]}\|=1$, these operators are defined token-wisely as:
@@
where $\bm{\gamma}\in\mathbb{R}^{D}$ is a trainable parameter vector. RMSNorm rescales each token to have unit norm and applies element-wise scaling using the learned parameter $\bm{\gamma}$, while $\Pi$ can be interpretted as the special case with $\bm{\gamma}=\mathbf{1}$. As we will show in Section 5, the trainable parameter $\bm{\gamma}$ plays an important role in stabilizing Jacobians.
@@
where $\eta$ denotes a discrete step size. The Omega layer ($\operatorname{Omg}$) is given by a linear transformation by anti-symmetric matrices and determines the rotational dynamics of oscillators. The projection operator $\operatorname{Proj}_{\bm{X}}(\bm{Y})$ and the normalization operator $\Pi(\bm{Y})$ are applied independently to each oscillator. We use the notation $\bullet^{(\text{osc})}$ to denote oscillator-wise operations. We provide further details of AKOrN in Appendix C.
@@
Previous studies on looped architectures (Saunshi et al., 2025; Bansal et al., 2022) have shown that injecting the input $\bm{C}\in\mathbb{R}^{S\times D}$ into the loop is important for achieving test-time scaling. As we show later, our theory (Proposition 5.1) indicates that the normalization layer plays a critical role in controlling the norm of the Jacobian matrix, particularly in the case of RMS normalization, which is widely used in practice. Based on these insights, we propose and investigate the following update rule, referred to as iterative self-attention (ItrSA):
@@ 4 Energy-based analysis
Propositions 4.1 and 4.2 imply that certain structures of weight matrices are desirable to ensure the existence of an energy function. Specifically, $W_{Q}W_{K}^{\top}$ can be asymmetric, whereas $W_{V}$ should remain symmetric. In the multi-head scenario, a low-rank structure in the QK product is required, which is consistent with practical Transformers, as they typically assume a (similar but different) low-rank structure as well. We refer to architectures that incorporate these properties as generalized symmetric SA, and we will explore their effectiveness in our experiments (Section 6.2). The proofs are provided in Sections A.2 and A.3.
@@ 5 Jacobian-based analysis
In general, energy functions are used to guarantee the convergence of dynamics to fixed points (a.k.a. Lyapunov’s direct method) (Khalil, 2002). While this is a concrete approach to achieve stable dynamics, the construction of energy functions is usually unsystematic, and thus it is not obvious whether we can handle more realistic SA dynamics (e.g., discrete updates with normalization). Furthermore, recent experimental results have reported non-stationary dynamics (e.g., oscillations) (Karagodin et al., 2024; Miyato et al., 2025), suggesting the need for more flexible approaches applicable to richer dynamics. Thus, we turn to analyzing the Jacobian matrix of state updates. The Jacobian controls the Lipschitzness of the function, and also naturally appears in linear stability analysis (a.k.a. Lyapunov’s indirect method), where state updates are locally described by $\bm{f}(\bm{x}+\Delta\bm{x})\approx\bm{f}(\bm{x})+\bm{J}(\bm{x}_{t})\Delta\bm{x}$ with the Jacobian $\bm{J}\coloneqq\partial\bm{f}/\partial\bm{x}$.
5.1 Normalization of spectral norm
@@
$\left\|\frac{\partial\operatorname{RMSNorm}(\bm{X}+\eta\Delta\bm{X})}{\partial\bm{X}}\right\|_{2}$ $\leq\frac{\max_{j}(|\gamma_{j}|)}{R}\left(1+|\eta|\left\|\bm{J}_{\text{MSA}}(\bm{X})\right\|_{2}\right),$ (14)
@@
We show the proof in Section A.4. This proposition highlights the key stabilizing effect of normalization: the spectral norm is inversely proportional to $R$. This effect appears to be particularly significant for preventing signal explosion in looped architectures, where the same operation is repeatedly applied.
Figure 2 shows that, in a practical model, normalization reduces the maximum Lyapunov exponent. This exponent corresponds to a time-averaged maximum singular vale of $\bm{J}$, which is further explained and analyzed in Section 5.3. This result supports the stabilizing effect of normalization implied by Proposition 5.1.
Jacobian eigenvalues of SA. Castin et al. (2024) provide an upper bound on the maximum eigenvalue (in the form of a Lipschitz constant) of the Jacobian of SA defined as $\bm{J}_{\text{MSA}}(\bm{X})\coloneqq\partial\operatorname{MSA}(\bm{X})/\partial\bm{X}$. Specifically, their results in Theorem 3.3 and Lemma 3.8 state that the Jacobian $\bm{J}_{\text{MSA}}(\bm{X})$ satisfies
@@
for input tokens $\bm{X}\in\mathbb{R}^{S\times D}$ such that $\|\bm{X}_{[i,:]}\|\leq r$ for all $i\in[1,S]$. This inequality implies that when the norms of the weight matrices $\|\bm{W}_{h}^{O}\|_{2},\|\bm{W}^{V}_{h}\|_{2}$, or $\|\bm{W}^{Q}_{h}\bm{W}^{K\top}_{h}\|_{2}$, or the number of tokens $S$ becomes large, the spectral norm can also become large.
@@
This is precisely the issue that normalization techniques can address. Figure 1 demonstrates that the spectral norm of the untrained SA’s Jacobian can be effectively reduced through normalization (see Section B.4 for experimental settings). Interestingly, we empirically observed that the spectral norm is not only reduced by normalization but remains $O(1)$ to the number of tokens. This suggests that the current theoretical bound is conservative, and a tighter bound remains future theoretical work.
@@ 5.2 Normalization of oscillatory components
$\dot{\bm{x}}$ $=\bm{\Omega}\bm{x},$ $\bm{J}(\bm{x})$ $=\bm{\Omega}$ (continuous) (16)
$\bm{x}^{(t+1)}$ $=(\bm{I}_{D}+\eta\bm{\Omega})\bm{x}^{(t)},$ $\bm{J}(\bm{x}^{(t)})$ $=\bm{I}_{D}+\eta\bm{\Omega}$ (discrete w/o Norm) (17)
$\bm{x}^{(t+1)}$ $=\Pi\big{(}(\bm{I}_{D}+\eta\bm{\Omega})\bm{x}^{(t)}\big{)},$ $\bm{J}(\bm{x}^{(t)})$ $=\left(\bm{I}_{D}-\frac{\bm{y}\bm{y}^{\top}}{\|\bm{y}\|^{2}}\right)\frac{\bm{I}_{D}+\eta\bm{\Omega}}{\|\bm{y}\|}$ (discrete w/ Norm) (18)
@@
It is known that the pure imaginary eigenvalues in the continuous-time limit are essential for capturing long-term signal dependencies, but can be significantly damaged by discretization (Chang et al., 2019). Generally, in continuous-time systems $\dot{\bm{x}}=\bm{f}(\bm{x})$, the equilibrium point is Lyapunov stable if all eigenvalues $\lambda_{j}$ of the Jacobian $\bm{J}(\bm{x})$ satisfy $\operatorname{Re}(\lambda_{j})\leq 0$. In Eq. (16), the Jacobian $\bm{J}(\bm{x})=\eta\bm{\Omega}$ has purely imaginary eigenvalues and they are on the boundary of stability, allowing persistent oscillations. In contrast, the equilibrium points of discrete-time systems $\bm{x}^{(t+1)}=\bm{f}(\bm{x}^{(t)})$ are Lyapunov stable if all eigenvalues $\lambda_{j}$ of the Jacobian satisfy $|\lambda_{j}|\leq 1$. For Eq. (17), all eigenvalues of the Jacobian $\bm{J}(\bm{x})=\bm{I}_{D}+\eta\bm{\Omega}$ take the form $1\pm i\eta\omega_{j}$ for $\omega_{j}\geq 0$, implying that $|\lambda_{j}|\geq 1$. Therefore, the system becomes unstable. To avoid the fundamental instability arising from discretization, previous work on architecture design inspired by dynamical systems proposed to add a diffusion term to the anti-symmetric weight matrix, i.e., $\bm{\Omega}-\gamma\bm{I}$ ($\gamma>0$) (Haber and Ruthotto, 2017; Chang et al., 2019).
We find that a normalization layer (18) serves as an alternative way to mitigate this instability by effectively rescaling the system through division by $\|\bm{y}\|$. For simplicity, suppose that all eigenvalues of $\Omega$ degenerate to the same $\omega_{j}=\omega$. After a straightforward calculation, we obtain $|\lambda_{j}|\leq 1$ (see Section A.5 for the derivation). The effect of this normalization is illustrated in Figure 2. Although this scenario represents an idealized setting, the normalization of imaginary components is empirically observed even in the case of SA, as shown in Figure 2.

### v3

@@ Abstract
The theoretical understanding of self-attention (SA) has been steadily progressing. A prominent line of work studies a class of SA layers that admit an energy function decreased by state updates. While it provides valuable insights into inherent biases in signal propagation, it often relies on idealized assumptions or additional constraints not necessarily present in standard SA. Thus, to broaden our understanding, this work aims to relax these energy constraints and provide an energy-agnostic characterization of inference dynamics by dynamical systems analysis. In more detail, we first consider relaxing the symmetry and single-head constraints traditionally required in energy-based formulations. Next, we show that analyzing the Jacobian matrix of the state is highly valuable when investigating more general SA architectures without necessarily admitting an energy function. It reveals that the normalization layer plays an essential role in suppressing the Lipschitzness of SA and the Jacobian’s complex eigenvalues, which correspond to the oscillatory components of the dynamics. In addition, the Lyapunov exponents computed from the Jacobians demonstrate that the normalized dynamics lie close to a critical state, and this criticality serves as a strong indicator of high inference performance. Furthermore, the Jacobian perspective also enables us to develop regularization methods for training and a pseudo-energy for monitoring inference dynamics.
@@ 1 Introduction
The theoretical understanding of self-attention (SA), a central component of Transformer architectures (Vaswani et al., 2017), has deepened in recent years, including phenomena such as rank collapse (Noci et al., 2022) and expressive capacity (Yun et al., 2020; Kajitsuka and Sato, 2024; Kingma and Ba, 2025). One major line of research formulates attention mechanisms as processes that minimize explicit or implicit energy functions (Ramsauer et al., 2021; Yang et al., 2022; Hoover et al., 2023). Since these energy functions serve as potential functions for gradient flows or as Lyapunov functions, they offer convergence guarantees and provide intuitive explanations for behaviors such as clustering and rank collapse in recurrent SA architectures (Geshkovski et al., 2025, 2023, 2024; Bruno et al., 2025). They restrict attention dynamics to a hypersphere, typically as a result of normalization. This facilitates theoretical analysis and can yield well-behaved functional properties (Castin et al., 2024).
@@
  • First, we revisit the energy-based formulation and partially relax traditional architectural constraints, such as symmetric weights and single-head assumptions, to better approximate realistic SA settings (Section˜4). These relaxed constraints provide insights into designing regularization methods, which we experimentally explore later in Section˜6.2.
  • To study a broader class of SAs more flexibly, we next analyze the Jacobian matrix of SA dynamics (Section˜5). The Jacobian approach is more general than the energy-based analysis (a.k.a. Lyapunov’s direct method) in the sense that it characterizes linear stability (a.k.a. Lyapunov’s indirect method) and enables us to more easily detect non-stationary dynamics, including oscillations. We demonstrate that normalization layers, unique to discrete updates, play a critical role in stabilizing dynamics. Specifically, they effectively suppress the Jacobian’s spectral norm (Proposition˜5.1) and control oscillatory behaviors by normalizing the complex eigenvalues of the Jacobian (Section˜5.2). In addition to the understanding of the normalization role, we empirically reveal that high-performance SA models exhibit a maximum Lyapunov exponent close to zero, suggesting that rich non-stationary inference dynamics emerge at the boundary between convergence and instability.
  • Finally, we investigate test-time scaling (performance improvement as the number of iterations increases) through the lens of Jacobians. We show that regularizing the spectral norm of weight matrices in SA improves performance (Section˜6.2), and that the Jacobian offers an interpretation of the pseudo-energy proposed in prior work, linking it to large eigenvalues (Section˜6.3).
Thus, our work broadens the dynamical understanding of SA and highlights the usefulness of the Jacobians and the Lyapunov exponent as promising and fundamental tools for further exploration of realistic SA architectures.
@@ 2 Related work
Geshkovski et al. (2025, 2023, 2024); Karagodin et al. (2024); Bruno et al. (2025) formulated recurrent SA dynamics as interactions among tokens (“particles”), enabling theoretical analysis of phenomena such as meta-stable clustering and rank-one collapse. Their continuous-time dynamics monotonically decrease an energy (Lyapunov) function, typically requiring constraints like single-head attention, hyperspherical token states, and symmetric weights. Yang et al. (2022) similarly interpreted Transformers as alternating minimization of energy functions, though with stricter conditions on step sizes and fixed-point proximity. Ramsauer et al. (2021) formalized the attention mechanism as modern Hopfield networks, and Hoover et al. (2023); Hu et al. (2025) further developed energy functions for Transformers. We do not address approaches based on the Hopfield networks, as they require architectural modifications distinct from standard self-attention.
@@ 3 Preliminaries
Geshkovski et al. (2025); Karagodin et al. (2024) used continuous equations and particle interpretation of tokens to model state-update dynamics of SA as:
@@
depending on the analyses. (Bruno et al., 2025) further assumes an unnormalized version of softmax. Under these conditions, the SA update can decrease an energy function. That is, the dynamics evolve in a way that monotonically decreases the energy, thereby ensuring the Lyapunov stability. Because these models suppose symmetric weights, we refer to a class of SA layers with symmetric weights and Lyapunov functions as symmetric SA. Yang et al. (2022) also formalized updates of SA using a symmetric matrix  (Appendix˜C).
@@
To facilitate theoretical analysis of SA, several studies (Geshkovski et al., 2025; Miyato et al., 2025) have introduced a spherical constraint on token vectors by enforcing that each token vector has unit norm. This constraint enables an interpretation of token interactions as dynamics of particles on a hypersphere, and plays a key role in controlling the Lipschitz continuity of SA (Castin et al., 2024). There are two commonly used operators with a spherical constraint: a normalization operator $\Pi$ that enforces the spherical constraint, and a projection operator $\operatorname{Proj}$ that projects onto the tangent space of the sphere. Given a token matrix $\bm{X},\bm{Y}\in\mathbb{R}^{S\times D}$ such that $\|\bm{X}_{[i,:]}\|=1$, these operators are defined token-wisely as:
@@
where $\bm{\gamma}\in\mathbb{R}^{D}$ is a trainable parameter vector. RMSNorm rescales each token to have unit norm and applies element-wise scaling using the learned parameter $\bm{\gamma}$, while $\Pi$ can be interpreted as the special case with $\bm{\gamma}=\mathbf{1}$. As we will show in Section˜5, the trainable parameter $\bm{\gamma}$ plays an important role in stabilizing Jacobians.
@@
where $\eta$ denotes a positive discrete step size. The Omega layer ($\operatorname{Omg}$) is given by a linear transformation by anti-symmetric matrices and determines the rotational dynamics of oscillators. The projection operator $\operatorname{Proj}_{\bm{X}}(\bm{Y})$ and the normalization operator $\Pi(\bm{Y})$ are applied independently to each oscillator. We use the notation $\bullet^{(\text{osc})}$ to denote oscillator-wise operations. We provide further details of AKOrN in Appendix˜C.
@@
Previous studies on looped architectures (Saunshi et al., 2025; Bansal et al., 2022) have shown that injecting the input $\bm{C}\in\mathbb{R}^{S\times D}$ into the loop is important for achieving test-time scaling. As we show later, our theory (Proposition˜5.1) indicates that the normalization layer plays a critical role in controlling the norm of the Jacobian matrix, particularly in the case of RMS normalization, which is widely used in practice. Based on these insights, we propose and investigate the following update rule, referred to as iterative self-attention (ItrSA):
@@ 4 Energy-based analysis
Propositions˜4.1 and 4.2 imply that certain structures of weight matrices are desirable to ensure the existence of an energy function. Specifically, $W^{Q}_{h}W^{K\top}_{h}$ can be asymmetric, whereas $W^{V}_{h}$ should remain symmetric. In the multi-head scenario, a low-rank structure in the QK product is required. This aligns with practical Transformers, as they typically exhibit a low-rank structure due to the small inner dimension (the width of $W^{Q}_{h},W^{K}_{h}$). We refer to architectures that incorporate these properties as generalized symmetric SA, and we will explore their effectiveness in our experiments (Section˜6.2). The proofs are provided in Sections˜A.2 and A.3.
@@ 5 Jacobian-based analysis
In general, energy functions are used to guarantee the convergence of dynamics to fixed points (a.k.a. Lyapunov’s direct method) (Khalil, 2002). While this is a concrete approach to achieving stable dynamics, the construction of energy functions is usually unsystematic, and thus it is not obvious whether we can handle more realistic SA dynamics (e.g., discrete updates with normalization). Furthermore, recent experimental results have reported non-stationary dynamics (e.g., oscillations) (Karagodin et al., 2024; Miyato et al., 2025), suggesting the need for more flexible approaches applicable to richer dynamics. Thus, we turn to analyzing the Jacobian matrix of state updates. The Jacobian controls the Lipschitzness of the function and also naturally appears in linear stability analysis (a.k.a. Lyapunov’s indirect method), where state updates are locally described by $\bm{f}(\bm{x}+\Delta\bm{x})\approx\bm{f}(\bm{x})+\bm{J}(\bm{x}_{t})\Delta\bm{x}$ with the Jacobian $\bm{J}\coloneqq\partial\bm{f}/\partial\bm{x}$.
5.1 Normalization and spectral norm
@@
$\left\|\frac{\partial\operatorname{RMSNorm}(\bm{X}+\eta\Delta\bm{X})}{\partial\bm{X}}\right\|_{2}$ $\leq\frac{\max_{j}(|\gamma_{j}|)}{R}\left(1+\eta\left\|\bm{J}_{\text{MSA}}(\bm{X})\right\|_{2}\right),$ (14)
@@
We show the proof in Section˜A.4. This proposition highlights the key stabilizing effect of normalization: the spectral norm is inversely proportional to $R$. This effect appears to be particularly significant for preventing signal explosion in looped architectures, where the same operation is repeatedly applied.
Figure˜2 shows that, in a practical model, normalization reduces the maximum Lyapunov exponent. This exponent corresponds to a time-averaged maximum singular value of $\bm{J}$, which is further explained and analyzed in Section˜5.3. This result supports the stabilizing effect of normalization implied by Proposition˜5.1.
In addition, under the assumption that $\|\Delta\bm{X}_{[i,:]}\|\geq\varepsilon$ for some constant $\varepsilon>0$ for all $i$, we can further show that as $\eta\to\infty$,
$\left\|\frac{\partial\operatorname{RMSNorm}(\bm{X}+\eta\Delta\bm{X})}{\partial\bm{X}}\right\|_{2}=O(1).$ (15)
That is, the Jacobian norm remains bounded even for a large $\eta$ (see Section˜A.4).
Jacobian eigenvalues of SA. Castin et al. (2024) provide an upper bound on the maximum eigenvalue (in the form of a Lipschitz constant) of the Jacobian of SA defined as $\bm{J}_{\text{MSA}}(\bm{X})\coloneqq\partial\operatorname{MSA}(\bm{X})/\partial\bm{X}$. Specifically, their results in Theorem 3.3 and Lemma 3.8 state that the Jacobian $\bm{J}_{\text{MSA}}(\bm{X})$ satisfies for input tokens $\bm{X}\in\mathbb{R}^{S\times D}$ such that $\|\bm{X}_{[i,:]}\|\leq r$ for all $i\in[1,S]$.
@@
This inequality indicates that the right-hand side of Eq. (14) is bounded independently of the input $\bm{X}$. It further implies that when the norms of the weight matrices $\|\bm{W}_{h}^{O}\|_{2},\|\bm{W}^{V}_{h}\|_{2}$, or $\|\bm{W}^{Q}_{h}\bm{W}^{K\top}_{h}\|_{2}$, or the number of tokens $S$ becomes large, the spectral norm can also become large. This is precisely the issue that normalization techniques can address. Figure˜1 demonstrates that the spectral norm of the untrained SA’s Jacobian can be effectively reduced through normalization (see Section˜B.4 for experimental settings). Interestingly, we empirically observed that the spectral norm is not only reduced by normalization but remains $O(1)$ with respect to the number of tokens. This suggests that the current theoretical bound (Eqs. (14) and (16)) is conservative, and a tighter bound remains future theoretical work.
@@ 5.2 Normalization of oscillatory components
$\dot{\bm{x}}$ $=\bm{\Omega}\bm{x},$ $\bm{J}(\bm{x})$ $=\bm{\Omega}\hskip-5.0pt$ (continuous) (17)
$\bm{x}^{(t+1)}$ $=(\bm{I}_{D}+\eta\bm{\Omega})\bm{x}^{(t)},$ $\bm{J}(\bm{x}^{(t)})$ $=\bm{I}_{D}+\eta\bm{\Omega}\hskip-5.0pt$ (discrete w/o Norm) (18)
$\bm{x}^{(t+1)}$ $=\Pi\big((\bm{I}_{D}+\eta\bm{\Omega})\bm{x}^{(t)}\big),$ $\bm{J}(\bm{x}^{(t)})$ $=\left(\bm{I}_{D}-\frac{\bm{y}\bm{y}^{\top}}{\|\bm{y}\|^{2}}\right)\frac{\bm{I}_{D}+\eta\bm{\Omega}}{\|\bm{y}\|}\hskip-5.0pt$ (discrete w/ Norm) (19)
@@
It is known that the pure imaginary eigenvalues in the continuous-time limit are essential for capturing long-term signal dependencies, but can be significantly damaged by discretization (Chang et al., 2019). Generally, in continuous-time systems $\dot{\bm{x}}=\bm{f}(\bm{x})$, the equilibrium point is Lyapunov stable if all eigenvalues $\lambda_{j}$ of the Jacobian $\bm{J}(\bm{x})$ satisfy $\operatorname{Re}(\lambda_{j})\leq 0$. In Eq. (17), the Jacobian $\bm{J}(\bm{x})=\bm{\Omega}$ has purely imaginary eigenvalues and they are on the boundary of stability, allowing persistent oscillations. In contrast, the equilibrium points of discrete-time systems $\bm{x}^{(t+1)}=\bm{f}(\bm{x}^{(t)})$ are Lyapunov stable if all eigenvalues $\lambda_{j}$ of the Jacobian satisfy $|\lambda_{j}|\leq 1$. For Eq. (18), all eigenvalues of the Jacobian $\bm{J}(\bm{x})=\bm{I}_{D}+\eta\bm{\Omega}$ take the form $1\pm i\eta\omega_{j}$ for $\omega_{j}\geq 0$, implying that $|\lambda_{j}|\geq 1$. Therefore, the system becomes unstable. To avoid the fundamental instability arising from discretization, previous work on architecture design inspired by dynamical systems proposed to add a diffusion term to the anti-symmetric weight matrix, i.e., $\bm{\Omega}-\gamma\bm{I}$ ($\gamma>0$) (Haber and Ruthotto, 2017; Chang et al., 2019).
We find that a normalization layer (19) serves as an alternative way to mitigate this instability by effectively rescaling the system through division by $\|\bm{y}\|$. For simplicity, suppose that all eigenvalues of $\Omega$ degenerate to the same $\omega_{j}=\omega$. After a straightforward calculation, we obtain $|\lambda_{j}|\leq 1$ (see Section˜A.5 for the derivation). The effect of this normalization is illustrated in Figure˜2. Although this scenario represents an idealized setting, the normalization of imaginary components is empirically observed even in the case of SA, as shown in Figure˜2.